OpenStudy (anonymous):

What is the value of the expression w – 4 when w = –17? my answer is -21

OpenStudy (igreen):

w - 4 when 2 = -17 -17 - 4 -17 - 4 = -21 You're correct!

OpenStudy (igreen):

when w = -17*

OpenStudy (anonymous):

kk thought so but wanted to double check